OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = one halfx - 8. y = one halfx + 7 y = one halfx + 5 y = 2x + 7 y = 2x + 5

OpenStudy (tkhunny):

Have you considered the point-slope form? Build it!!

OpenStudy (anonymous):

Im confused can you explain

OpenStudy (tkhunny):

If I told you to build the equation of a line, given that it passes through (-2,3) and has a slope of 1/2, could you do it?

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

Okay, do it. Let's see what you get.

OpenStudy (anonymous):

Can you give me an example because im stuck

OpenStudy (tkhunny):

I was waiting for your example. You said you could do it. Am I doing this one for you?

OpenStudy (anonymous):

Nope but it would be helpful if you could show me an example so I can kinda like figure it out

OpenStudy (tkhunny):

Point-Slope Form Point:(h,k) Slope: m (y-k) = m(x-h) In your case: Point: (h,k) = (-2,6) Slope: m = 1/2 (y-6) = (1/2)(x+2) That's really it, but it wants it solved for y. y = (1/2)x + 1 + 6 = (1/2)x + 7 -- Looks like we have a winner!

OpenStudy (anonymous):

Ty

OpenStudy (tkhunny):

Let's see you do one.